OpenStudy (lisa123):

f(x)=x^2+1, g(x)=x-2 (f+g)(t-8)

OpenStudy (anonymous):

is there another function for t?

OpenStudy (lisa123):

no

OpenStudy (lisa123):

thats why im confused

OpenStudy (anonymous):

\[f(x)=x^2+1\\f(t-8)=(t-8)^2+1\] clean up with some algebra \[g(x)=x-2\\ g(t-8)=t-8-2\] clean up with some arithmetic then for \(f+g\) put a plus sign between them

OpenStudy (lisa123):

Thank you @satellite73 I got it right!